OpenStudy (anonymous):

What is the y−coordinate of the y−intercept of the line that passes through the points (−4, −4) and (4, 8)?

OpenStudy (anonymous):

ok so the best thing we can do is to find the equation of this line

OpenStudy (anonymous):

can you find the slope of this line ?

OpenStudy (anonymous):

how do you find it again??

OpenStudy (anonymous):

i havent done this is so long..:/

OpenStudy (anonymous):

in*

OpenStudy (anonymous):

ok so slope : m = (y1 - y2) / (x1-x2)

OpenStudy (anonymous):

oh oh ok let me try

OpenStudy (anonymous):

-4-8=-12 and -4-4=-8 so -12over-8?

OpenStudy (anonymous):

yes .. which is 1.5

OpenStudy (anonymous):

now.. find line equation by : y = y1 + m(x-x1) (x1,y1) may be one of your two points that you got

OpenStudy (anonymous):

wait what? :P

OpenStudy (anonymous):

y = y1 + m(x-x1) -> if you use this formula you can get the line equation in our case we can choose x1 = -4 , y1 = -4 or x1= 4 ,y1 = 8

OpenStudy (anonymous):

so the anwser is 4,8?

OpenStudy (anonymous):

or wait? yah no...im lost..im sorry :(

OpenStudy (anonymous):

ok again .. we need to find y intercept

OpenStudy (anonymous):

its the point where x=0 right ?

OpenStudy (anonymous):

i thnk soo yes

OpenStudy (anonymous):

so what i want to do is to find the equation of the line what it means is that in this equation i will be able to plug x =0 and get the y-intercept or any other x that i want to find the corresponding y of it

OpenStudy (anonymous):

we already found the slop of this line m = 1.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now if you have a point that the line passes through and you got also the slope of the line you could find the equation that describes the line using this formula : y = y1 + m(x-x1) where m - slope (x1,y1) a point that the line passes through

OpenStudy (anonymous):

oh ok so it could be y=8+1.5(4+4) ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why ? look at the formula 1.5(x-x1) so y=8+1.5(x-4)

OpenStudy (anonymous):

do you understand why ? x for us is anything we want in this case we would like to plug x=0 since we want y intercept

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

now y=8+1.5(x-4) this is our line equation replace x by 0 and you will get y intercept .. do you understand what we have done ?

OpenStudy (anonymous):

yes i think so..so y=8+1.5(0)-6 ?

OpenStudy (anonymous):

yes y intercept would be 2

OpenStudy (anonymous):

ohhhhhh thats the anwser!! :) thank you! can you help me with a few more??

OpenStudy (anonymous):

yes but you must understand .. try to do it by your self and tell me where you stuck

OpenStudy (anonymous):

okay the questions i put up here i dont understand..:P let me tryyy